OpenStudy (anonymous):

what are the x-intercepts for the graph of y=3(x-1)(x+6)? A)(1,0) and (-6,0) B)(0,-1) and (0,6) C)(0,1) and (0,-6) D)(-1,0) and (6,0)

OpenStudy (ahsome):

Your equation is: \(y=3(x-1)(x+6)\) Firstly, expand the brackets

OpenStudy (anonymous):

what is it?

OpenStudy (ahsome):

Do you know how to expand brackets?

OpenStudy (anonymous):

no

OpenStudy (ahsome):

Have you heard of the foil method before?

OpenStudy (anonymous):

what is the answer to this one?

OpenStudy (ahsome):

I can't tell you the answer, only help you

OpenStudy (ahsome):

Hello?

OpenStudy (anonymous):

I don't get it

OpenStudy (ahsome):

Do you know how to expand a bracket?

OpenStudy (anonymous):

nope

Directrix (directrix):

> y=3(x-1)(x+6) The x-intercepts lie on the x-axis and have a y-value of 0. So, you want to solve 0 = 3(x-1)(x+6) Divide both sides by 3. 0 = (x-1)(x+6)

Directrix (directrix):

Apply the Zero Product Property to get: 0 = (x-1)(x+6) so 0 = x -1 or 0 = x + 6 x = ? or x = ? After you solve for x, be careful to place the x-intercepts in the proper position in the (x,y) format. (1,0) and (-6,0) is not the same as (0,1) and (0,-6).